OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of lemon juice are present in the bottle? Show your work. (5 points) Please Help I Will Give Medal .

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

what have you already done? can you show?

OpenStudy (anonymous):

I Came Up With A equation t=60t +40t :(

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

hmm i don't know if i did it correct but i'll try to explain i made 2 sides to the equation left side - is where we make equation to find for X- which is how many liters of Lemon juice is there and right side - we sum up X with remaining 1 liter of water

OpenStudy (anonymous):

Okay So How Would It Look @Somy :)

OpenStudy (somy):

okay im not too sure but i made it this way 40X - 60 = 1 + X here we find X for liters of lemon juice

OpenStudy (anonymous):

Okay :) They Make These Questions So Hard :( Its Frustrating

OpenStudy (somy):

but that i just used math logic depending on my final wanted answer coz i want X to be equal to = 60/40 coz it leads me directly to answer for liters of juice and the other side i just decided to do addition part to give the total liters as they wanted

OpenStudy (somy):

would be good if math ppl cross check tho :P coz math is not one of my strongest points kinda

OpenStudy (anonymous):

Okay && Me Either Smh I Struggle Everyday :( @Somy Thanks Fro Trying

OpenStudy (somy):

no problem :) and good luck!

OpenStudy (anonymous):

Thanks :)